OpenStudy (anonymous):

HELP PLEASE! Sam is putting money into a savings account. He starts with $550 in the savings account, and each week he adds $40.Let S represent the total amount of money in the savings account (in dollars), and let W represent the number of weeks Sam has been adding money. Write an equation relating S to W. Then use this equation to find the total amount of money in the savings account after 18 weeks.

OpenStudy (danjs):

one point on the line (0,550) Slope of the line: 40

sammixboo (sammixboo):

Well have you written an equation yet?

OpenStudy (anonymous):

HUH? @DanJS i need an equation

OpenStudy (danjs):

you have one point and the slope, you can use point slope form for a line S - S1 = m(W-W1) Where m = 40, and (W1,S1) = (0,550)

OpenStudy (danjs):

plugging the numbers in gives S - 550 = 40(x - 0)

OpenStudy (anonymous):

i dont need a line i need an equation

OpenStudy (danjs):

S- 550 = 40*W S = 40*W + 550

OpenStudy (anonymous):

and i need the amount of money in 18 weeks

OpenStudy (danjs):

the equation is a line

OpenStudy (danjs):

so when W= 18 weeks, calculate S S = 40*W+550

OpenStudy (anonymous):

ohhhh so the equation is S - S1 = m(W-W1)???

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

yes, that is the point-slope form of a line if you are given a point (S1,W1) and the slope m, you can plug the values in and get the equation for the line

OpenStudy (anonymous):

okay got it

OpenStudy (anonymous):

so now how much money would it be in 18 weeeks?

OpenStudy (danjs):

Let W=18

OpenStudy (danjs):

S = 40*18+550

OpenStudy (danjs):

1270

OpenStudy (danjs):

the point (18,1270) is also on the line

OpenStudy (anonymous):

@DanJS can i write the equation without brackets??

OpenStudy (danjs):



OpenStudy (danjs):

yeah from point slope form you get S - 550 = 40(W - 0) Rearranging some, you get S = 40*W + 550

OpenStudy (anonymous):

@DanJS hey

OpenStudy (danjs):

?

OpenStudy (anonymous):

so the money in 18 weeks would be 1270 right

OpenStudy (danjs):

yep

OpenStudy (danjs):

each of these values , 0 weeks 550 dollars, and 18 weeks 1270 dollars , are just points on a line (S,W) -- (0 , 18) and (18,1270) S = 40W+550